Shown is the boy riding on a cart, throwing the ball horizontally with speed v1 = 5 m / s. Find the speed of the cart after throwing the ball, when it is known that the mass of the ball is m1 = 500 g, while the mass of the boy together with the mass of the cart is m2 = 70 kg. At first the cart was at rest. The frictional forces between the cart and the ground are not taken into account


1
Expert's answer
2021-01-28T09:57:28-0500

By the law of conservation of momentum, we have:


"p_i=p_f,""m_1v_1+m_2v_2=0,""v_2=-\dfrac{m_1v_1}{m_2},""v_2=-\dfrac{0.5\ kg\cdot5\ \dfrac{m}{s}}{70\ kg}=-0.036\ \dfrac{m}{s}."

The sign minus means that the speed of the cart after throwing the ball directed in the opposite direction to the motion of the ball.